OpenStudy (anonymous):

The length of a rectangle is 2 feet more than twice the width. The perimeter is 34 feet. Find the dimensions of the rectangle.

OpenStudy (anonymous):

if the width is x the length will be 2x+2. the perimeter is 2(width + length) 2(x + 2x + 2) = 34 2(3x +2) = 34 6x + 4 = 34 6x = 30 x = 5

OpenStudy (anonymous):

so the width is 5 and the length will be 12

OpenStudy (anonymous):

Thank you so much!!!! This really helped :)